An airplane accelerates down a runway at 3.20 m/s2 for 32.8 s until is finally lifts off the ground. Determine the distance traveled before takeoff.


1
Expert's answer
2021-09-18T14:59:42-0400
"s=v_0t+\dfrac{1}{2}at^2,""d=0\ \dfrac{m}{s}\cdot32.8\ s+\dfrac{1}{2}\cdot3.20\ \dfrac{m}{s^2}\cdot(32.8\ s)^2=1721\ m."